OpenStudy (anonymous):

Jessica spun this spinner 30 times with the following results. According to the data, what is the probability of the spinner landing on red twice in two spins? Red 10 Yellow 15 Green 2 Blue 3 spinner has 2 red, 2 yellow, 1 green and 1blue

OpenStudy (turingtest):

total colors spaces=2+2+1+1=6 2 are red so odds of hitting red are 2/6=1/3 twice in a row the odds are 1/3*1/3=1/9. I don't see what the data has to do with the odds.

OpenStudy (anonymous):

i dont either it included it but i think it's just there to confuse me